# From Prompts to Protection: Large Language Model-Enabled In-Context Learning for Smart Public Safety UAV

### v1

@@ IV-C Data Collection Schedule
As depicted in Fig. 3, the UAV hovers over the disaster area, enabling it to approach ground sensors closely and utilize short-range LoS communication links for efficient data collection. However, selecting a specific ground sensor for data collection may result in buffer overflows at other ground sensors if their data queues are near capacity while new data continues to accumulate. Moreover, ground sensors located farther from the UAV typically experience degraded channel conditions, increasing the likelihood of transmission errors. Therefore, an effective data collection schedule is essential to mitigate data queue overflows and reduce communication failures[13]. LLM-enabled ICL is crucial for data collection scheduling in UASNETs, as it facilitates dynamic adaptation to changing environments. By leveraging ICL, the UAV can continuously adjust its data collection scheduling strategies in real- time, selecting the most appropriate sensors. This enables the system to make informed, context-aware decisions without retraining, thereby optimizing data collection efficiency and responsiveness in dynamic or mission-critical scenarios. This ensures efficient, reliable, and autonomous data collection operations, enhancing overall system performance. The relevant task description is depicted in 2(e). This system uses an LLM to optimize data collection scheduling for minimal packet loss by processing real-time sensor inputs (queue lengths and channel conditions) while maintaining constraints that prioritize sensors with high queue lengths (to avoid buffer overflows) and those with good channel conditions (to ensure reliable transmission). The LLM generates an adaptive collection schedule, which is then deployed, recording the resulting system performance (packet loss) and the latest environmental state as feedback. This feedback loop, consisting of the LLM’s decisions, the latest system state, and performance metric, enables continuous improvement by providing information for subsequent planning decisions, allowing the LLM to iteratively refine its policy to minimize packet loss.
@@ IV-D UAV Power Control
We need to optimize energy efficiency and power control in UASNETs because the UAV and ground sensors have limited energy resources, making the optimization of energy and power consumption crucial for the overall performance of the UASNETs. DRL solutions have been proposed to exploit the network’s resources in a near-optimal manner, aiming to maximize energy efficiency effectively. However, these solutions require complex training, and their deployment is not rapid and straightforward. LLM-enabled ICL can be leveraged to optimize the transmit power in UASNETs to maximize overall energy efficiency.
Task description for UAV power control is depicted in 2(c) with a task objective, and the example set represents operational state and environmental context to assist the LLM in inferring optimal power control strategies. This system leverages an LLM to dynamically optimize UAV transmit power for maximum energy efficiency by processing real-time power levels and channel conditions, while adhering to constraints prioritizing sensors with low remaining energy (preventing node failure) and favoring those with strong channel conditions (ensuring reliable transmission). The LLM generates power adjustment commands executed, with the resulting system performance and updated operational state being captured as feedback. This closed-loop mechanism, incorporating the LLM’s power decisions, the latest UAV status, and communication outcomes, enables continuous refinement of power allocation strategies, progressively allowing the UAV to achieve optimal energy efficiency. This approach enables adaptive decision-making without the need for complex training. Such optimization is particularly valuable in public safety scenarios such as emergency response, where efficient power usage directly impacts mission duration and system performance.
(a) Normal and Under attack operation of ICLDC where the jailbreaking attack happens.
(b) Performance with changing the number of ground sensors.
Figure 4: Performance analysis of ICLDC in two scenarios with 10 ground sensors.
@@ IV-E Transmit Power of Ground Sensors
Power control in UASNETs, which involves explicitly managing the transmit power of ground sensors, plays a key role in enhancing system performance by optimizing interference management, signal quality, and resource usage, all while accounting for the limited battery capacities of the ground sensors. This system employs an LLM to optimize ground sensor transmit power for maximal energy efficiency by analyzing current power levels and channel conditions while following constraints that prioritize sensors with high queue lengths (to prevent buffer overflow) and favor good channel conditions. Based on the task description depicted in 2(d), the LLM generates power adjustment recommendations that are implemented, with the resulting system performance and last environmental state being monitored as feedback. This continuous feedback loop, tracking the LLM’s power decisions, the latest sensor data, and transmission outcomes, enables dynamic refinement of power settings, allowing the system to progressively improve energy efficiency while maintaining reliable data transmission and preventing queue overflows. LLM-enabled ICL offers a lightweight alternative to DRL for power control in UASNETs, where it replaces retraining with example-based reasoning and enables real-time decision-making with minimal computational overhead, which is especially well-suited for dynamic, resource-constrained, and safety-critical environments where DRL’s complexity and retraining demands become limiting factors.
@@
Training Requirements Complex model training Improved Model Training No retraining or parameter updates
Deployment Speed Not compatible with urgent requirements Improves real-time performance Rapid adaptability and deployment
@@
Key Advantage State-of-the-art for UASNETs Robust solution for complex scenarios Flexible and generalizable
Key Limitation Slow training and deployment Still requires complex integration Convergence properties unproven
V Case Study: Data Collection Schedule
This section presents a case study focused on data collection scheduling. The objective is to demonstrate how ICL effectively manages data collection scheduling and is vulnerable to jailbreaking attacks. An ICL-based Data Collection Scheduling (ICLDC) framework for UASNETs is proposed in [14]. The proposed ICLDC leverages ICL to optimize the UAV’s emergency data collection schedules. Specifically, the UAV gathers sensory information, such as queue lengths and channel conditions, from ground sensors and transmits this data to an edge-hosted LLM. The LLM is initialized with a structured system prompt that outlines its operational role, constraints, and ethical safeguards, ensuring compliance with safety and security requirements. Upon receiving the environmental data, the LLM generates a detailed task description in natural language, encompassing the mission objective, input parameters, operational rules, illustrative examples, expected outputs, and a feedback mechanism. Based on this context, the LLM infers an optimal data collection schedule, which the UAV subsequently executes. Performance metrics and updated environmental states are continuously monitored and incorporated as feedback into the evolving task description. This feedback is utilized in subsequent decision cycles, enabling the system to adapt dynamically and improve scheduling efficacy over time. Moreover, the paper examines the system’s vulnerability to adversarial manipulation, specifically, jailbreaking attacks, where the task description is maliciously altered to degrade scheduling performance and increase network cost. This highlights the critical need for robust safeguards when deploying LLM-enabled control schemes in mission-critical applications. Fig. 4(a) illustrates the performance of the proposed ICLDC under attack in comparison to the normal scenario. The results clearly show that the attack substantially degrades performance, resulting in higher costs. The attacker employs context-based manipulation to alter the task description within the LLM’s prompt. This deliberate interference misleads the LLM, prompting it to choose suboptimal or potentially harmful actions, which in turn disrupts the overall system behavior. Fig. 4(b) illustrates the performance of ICLDC compared to the under-attack baseline as the number $N$ of ground sensors increases. In general, a larger number of sensors leads to higher costs due to an increased likelihood of buffer overflow. Notably, ICLDC consistently achieves the lowest cost, demonstrating its effectiveness in managing network dynamics.
@@ VI Discussion: From DRL to LLM Optimizers
Recent advancements in prompting techniques have significantly enhanced the performance of LLMs across diverse domains. Leveraging their intrinsic ability to comprehend and generate natural language, LLMs introduce a novel paradigm for addressing optimization problems. Conventional optimization approaches typically require explicit formalization of the problem and the derivation of update rules through specialized solvers. In contrast, LLM-enabled optimization frameworks allow the problem to be expressed in natural language, enabling the model to iteratively propose candidate solutions based on the problem description and previously generated outputs.
@@
The LLM-enabled optimization framework offers several important advantages. First, it facilitates rapid adaptation to different UAV tasks by simply modifying the problem description contained in the prompt. Second, the iterative nature of the process enables continuous refinement of solutions without the need to retrain or change the underlying model parameters. Finally, the optimization process can be effectively customized through prompt engineering by including specific instructions that enforce desired solution properties such as feasibility constraints or performance criteria.
Such ICL capabilities enable LLMs to act as flexible and generalizable optimization agents. This change has the potential to accelerate the development of solutions in UASNETs where explicit mathematical modelling is a challenge. However, further research is needed to accurately quantify the convergence properties and solution quality that can be achieved by LLM-based optimization frameworks.
@@
Overall, while DRL algorithms are state of the art for optimizing UASNETs, they reach their limits in emergency scenarios due to their slow training, complex application and gaps between simulation and reality, making them unsuitable for time-critical situations. DMs improve DRL by generating synthetic data and improving the adaptability of policies, but are still computationally intensive. In contrast, LLM optimizers that use ICL offer a promising alternative. They enable fast, training-free adaptation through natural language prompts, iterative refinement of the solution, and optimization that can be interpreted by humans - important advantages for emergency response. However, LLM-based approaches require further validation of their convergence properties and real-world reliability, suggesting a potential for hybrid systems that combine the robustness of DRL with the flexibility of LLM to meet urgent public safety needs. Table II compares the Optimization Approaches for UASNETs in Emergency Scenarios
@@ VII Most recent trends
This section highlights the future direction of LLM-enabled ICL and the ways to improve the performance and deployment.
@@ VII-1 Enriching Examples using Diffusion Models
Focus on enhancing the generalization capabilities of public safety UAV by addressing the scarcity of high-quality training data. Develop methods to generate diverse and realistic synthetic examples using DMs to augment limited emergency datasets. Further research should explore optimizing DM-based data generation to ensure the quality and relevance of synthetic examples for robust generalization.
@@ VII-3 Enhanced Robustness Against Adversarial Attacks
Investigate and develop advanced defense mechanisms to mitigate vulnerabilities such as jailbreaking attacks, which manipulate prompts to bypass built-in safety mechanisms and cause LLMs to generate unauthorized or malicious content. Address both white-box and black-box categories of adversarial attacks to ensure the reliability of LLM-enabled ICL in mission-critical scenarios.
@@ VII-4 Extension to Public Safety UAV Swarms
Extend the application of LLM-enabled ICL to public safety UAV swarms, focusing on facilitating coordination among multiple UAVs. Develop methods that enable cooperative task execution in the swarm, leveraging the contextual understanding and adaptive reasoning capabilities of LLMs to improve overall swarm performance and mission success.
@@ VII-5 Real-World Validation and Transferability
Conduct extensive field trials to evaluate the performance of LLM-enabled ICL in real-world public safety applications. Explore the transferability of LLM-enabled ICL to other UAV applications, such as agriculture and infrastructure inspection, by repurposing task descriptions and prompts.

### v2

@@ IV-C Data Collection Schedule
As depicted in Fig. 3, the UAV hovers over the disaster area, enabling it to approach ground sensors closely and utilize short-range LoS communication links for efficient data collection. However, selecting a specific ground sensor for data collection may result in buffer overflows at other ground sensors if their data queues are near capacity while new data continues to accumulate. Moreover, ground sensors located farther from the UAV typically experience degraded channel conditions, increasing the likelihood of transmission errors. Therefore, an effective data collection schedule is essential to mitigate data queue overflows and reduce communication failures. LLM-assisted ICL is crucial for data collection scheduling in UASNETs, as it facilitates dynamic adaptation to changing environments. By leveraging ICL, the UAV can continuously adjust its data collection scheduling strategies in real- time, selecting the most appropriate sensors. This enables the system to make informed, context-aware decisions without retraining, thereby optimizing data collection efficiency and responsiveness in dynamic or mission-critical scenarios. This ensures efficient, reliable, and autonomous data collection operations, enhancing overall system performance. The relevant task description is depicted in 2(e). This system uses an LLM to optimize data collection scheduling for minimal packet loss by processing real-time sensor inputs (queue lengths and channel conditions) while maintaining constraints that prioritize sensors with high queue lengths (to avoid buffer overflows) and those with good channel conditions (to ensure reliable transmission). The LLM generates an adaptive collection schedule, which is then deployed, recording the resulting system performance (packet loss) and the latest environmental state as feedback. This feedback loop, consisting of the LLM’s decisions, the latest system state, and performance metric, enables continuous improvement by providing information for subsequent planning decisions, allowing the LLM to iteratively refine its policy to minimize packet loss. In this scheduling task, the cost is evaluated using two observable components: the average packet-loss rate across all active sensors and the backlog penalty reflecting the number of sensors with nearly full data queues. Both quantities are directly measurable through MAC-layer acknowledgments and queue-level reports and are summarized by the edge controller as contextual feedback to guide the next scheduling decision.
@@ IV-D UAV Power Control
We need to optimize energy efficiency and power control in UASNETs because the UAV and ground sensors have limited energy resources, making the optimization of energy and power consumption crucial for the overall performance of the UASNETs. DRL solutions have been proposed to exploit the network’s resources in a near-optimal manner, aiming to maximize energy efficiency effectively. However, these solutions require complex training. LLM-assisted ICL can be leveraged to optimize the transmit power and improve energy efficiency.
Task description for UAV power control is depicted in 2(c) with a task objective, and the example set represents operational state and environmental context to assist the LLM in inferring optimal power control strategies. This system leverages an LLM to dynamically optimize UAV transmit power for maximum energy efficiency by processing real-time power levels and channel conditions, while adhering to constraints prioritizing sensors with low remaining energy (preventing node failure) and favoring those with strong channel conditions (ensuring reliable transmission). The LLM generates power adjustment commands executed, with the resulting system performance and updated operational state being captured as feedback. This closed-loop mechanism, incorporating the LLM’s power decisions, the latest UAV status, and communication outcomes, enables continuous refinement of power allocation strategies, progressively allowing the UAV to achieve optimal energy efficiency. This approach enables adaptive decision-making without the need for complex training. Such optimization is particularly valuable in public safety scenarios such as emergency response, where efficient power usage directly impacts mission duration and system performance. For UAV-side power control, the cost is computed as the ratio between the total energy consumed for transmission and the number of successfully delivered bits during a given observation window. Additional penalties are introduced when the observed packet-loss rate or latency exceeds predefined quality-of-service thresholds. These values are measured by the UAV’s onboard power monitor and link-layer counters and then summarized as text feedback to the LLM for subsequent optimization.
@@ IV-E Transmit Power of Ground Sensors
Power control in UASNETs, which involves explicitly managing the transmit power of ground sensors, plays a key role in enhancing system performance. It optimizes interference management, signal quality, and resource usage, all while accounting for the limited battery capacities of the ground sensors. This system employs an LLM to optimize ground sensor transmit power for maximal energy efficiency. The LLM analyzes current power levels and channel conditions while adhering to constraints that prioritize sensors with high queue lengths (to prevent buffer overflow) and favor good channel conditions. Based on the task description depicted in 2(d), the LLM generates power adjustment recommendations that are implemented, with the resulting system performance and last environmental state being monitored as feedback. This continuous feedback loop, tracking the LLM’s power decisions, the latest sensor data, and transmission outcomes, enables dynamic refinement of power settings, allowing the system to progressively improve energy efficiency while maintaining reliable data transmission and preventing queue overflows. LLM-assisted ICL offers a lightweight alternative to DRL for power control in UASNETs, where it replaces retraining with example-based reasoning and enables real-time decision-making with minimal computational overhead, which is especially well-suited for dynamic, resource-constrained, and safety-critical environments where DRL’s complexity and retraining demands become limiting factors. In this sensor-side power control task, the cost reflects the overall energy efficiency of the network. It is derived from the total energy spent by all ground sensors divided by the number of successfully transmitted bits, together with an additional term representing the overall packet-loss ratio. These indicators are obtained directly from the sensors’ power telemetry and communication logs and are incorporated into the feedback context for subsequent LLM reasoning.
(a) Normal and under-attack operation of ICLDC.
(b) Performance with varying numbers of ground sensors.
(c) Measured LLM inference latency heatmap on UAV circular flight path.
Figure 4: Performance analysis of ICLDC in multiple scenarios with 10 ground sensors.
V Case Study: Data Collection Schedule
This section presents a case study focused on data collection scheduling. The objective is to demonstrate how ICL effectively manages data collection scheduling and is vulnerable to jailbreaking attacks. Our case study introduces a unique security vulnerability through its reliance on contextual prompts. To address this, a formal black-box attacker model is defined, where an adversary with no internal model access manipulates the system by injecting a malicious demonstration into the prompt context.
Table I: Runtime and Complexity Analysis of ICLDC
Metric ICLDC
Training Episodes 1 (30 steps)
Training Time $\sim$20 seconds (CPU)
Sample Efficiency 30 interactions
Compute Requirements Low (CPU-only)
Computational Complexity $\mathcal{O}(n^{2}+k\cdot n)$
The attacker’s objective is to alter the LLM’s output distribution to maximize a network-level loss function, such as one based on packet loss, thereby degrading system performance by forcing suboptimal decisions. We assume the attack is executed through methods such as jailbreaking via malicious prompts. An ICL-based Data Collection Scheduling (ICLDC) framework for UASNETs is proposed in [2]. The proposed ICLDC leverages ICL to optimize the UAV’s emergency data collection schedules. Specifically, the UAV gathers sensory information, such as queue lengths and channel conditions, from ground sensors and transmits this data to an edge-hosted LLM. The LLM is initialized with a structured system prompt that outlines its operational role, constraints, and ethical safeguards, ensuring compliance with safety and security requirements. Upon receiving the environmental data, the LLM generates a detailed task description in natural language, encompassing the mission objective, input parameters, operational rules, illustrative examples, expected outputs, and a feedback mechanism.
Based on this context, the LLM infers an optimal data collection schedule, which the UAV subsequently executes. Performance metrics and updated environmental states are continuously monitored and incorporated as feedback into the evolving task description. This feedback is utilized in subsequent decision cycles, enabling the system to adapt dynamically and improve scheduling efficacy over time.
Moreover, the paper examines the system’s vulnerability to adversarial manipulation, specifically jailbreaking attacks, where the task description is maliciously altered to degrade scheduling performance and increase network cost. This highlights the critical need for robust safeguards when deploying LLM-assisted control schemes in mission-critical applications. Fig. 4(a) illustrates the performance of the proposed ICLDC under attack in comparison to the normal scenario. The results clearly show that the attack substantially degrades performance, resulting in higher costs. The attacker employs context-based manipulation to alter the task description within the LLM’s prompt. This deliberate interference misleads the LLM, prompting it to choose suboptimal or potentially harmful actions, which in turn disrupts the overall system behavior. Fig. 4(b) illustrates the performance of ICLDC compared to the under-attack baseline as the number $N$ of ground sensors increases. In general, a larger number of sensors leads to higher costs due to an increased likelihood of buffer overflow. Notably, ICLDC consistently achieves the lowest cost, demonstrating its effectiveness in managing network dynamics. Fig. 4(c) shows a visualization of the LLM (GPT-4o-mini) inference latency measurements along a circular trajectory. The latency values at each point along the trajectory are represented by a color gradient, with different hues corresponding to different latency intensities measured in milliseconds (ms). The visualization effectively shows how LLM inference latency varies at different positions along the circular trajectory.
Table I shows that ICLDC requires only 1 training episode (30 steps), executes in approximately 20 seconds on a CPU, and achieves high sample efficiency (30 interactions), all while operating with low compute requirements. In addition, the computational complexity of ICLDC depends on both the length of the input sequence and the dimensionality of the model. Specifically, the cost increases with the square of the input sequence length, reflecting the overhead of processing interactions between elements in the sequence, and scales linearly with the model dimension for each input token. This characterization captures the computational cost of querying the language model at each decision step, providing a theoretical basis for evaluating the efficiency and scalability of the framework.
@@ V Case Study: Data Collection Schedule
Training Requirements Complex model training (1,000 episodes, 30,000 steps; $\sim$50 min on GPU) Improved Model Training No retraining or parameter updates (1 episode, 30 steps; $\sim$20 sec on CPU)
Deployment Speed Slow deployment Improves real-time performance Rapid adaptability and deployment
[…]
Compute Requirements High (GPU needed) Moderate Low (CPU-only; no GPU required)
@@
Inference Latency Low Low 0.3–0.7 ms (average 0.3 ms)
Key Advantage State-of-the-art for UAV-assisted data collection Robust solution for complex scenarios Flexible, generalizable, highly sample-efficient
Key Limitation Slow training and deployment Still requires complex integration Unproven convergence properties
@@ VI Discussion: From DRL to LLM Optimizers
Recent advancements in prompting techniques have significantly enhanced the performance of LLMs across diverse domains. Leveraging their intrinsic ability to comprehend and generate natural language, LLMs introduce a novel paradigm for addressing optimization problems. Conventional optimization approaches typically require explicit formalization of the problem and the derivation of update rules through specialized solvers. In contrast, LLM-assisted optimization frameworks allow the problem to be expressed in natural language, enabling the model to iteratively propose candidate solutions based on the problem description and previously generated outputs.
@@
The LLM-assisted optimization framework offers several important advantages. First, it facilitates rapid adaptation to different UAV tasks by simply modifying the problem description contained in the prompt. Second, the iterative nature of the process enables continuous refinement of solutions without the need to retrain or change the underlying model parameters. Finally, the optimization process can be effectively customized through prompt engineering by including specific instructions that enforce desired solution properties such as feasibility constraints or performance criteria.
Such ICL capabilities enable LLMs to act as flexible and generalizable optimization agents. This change has the potential to accelerate the development of solutions in UASNETs where explicit mathematical modeling is a challenge. However, further research is needed to accurately quantify the convergence properties and solution quality that can be achieved by LLM-based optimization frameworks. Specifically, LLM-assisted optimization remains challenging due to the autoregressive nature of inference and non-stationary context distributions. While we therefore avoid theoretical claims of convergence, the proposed framework does not rely on guaranteed convergence in practice. Instead, it operates under guarded conditions: i) Externally measured and bounded performance signals, such as packet loss, queue backlog and energy, are used as feedback; ii) The high-level action space is discretized and filtered by rule-based safety constraints; iii) The efficient optimization scheme avoids the complexity of dedicated model pre-training and fine-tuning, allowing for efficient correction and control. Together, these mechanisms provide operational stability and safety for edge-assisted UAV control.
Developing advanced memory architectures is essential to enhance the stability and contextual consistency of LLM-assisted control during extended missions. The current framework relies on feedback-driven iterative refinement, but its contextual understanding is limited to near-term interactions. For prolonged operations, maintaining coherent reasoning and avoiding repetitive strategies requires a structured representation of mission history.
@@
This can be addressed through a Conversation Memory Module comprising two complementary components: i) Short-Term Working Memory: A dynamic buffer that retains recent sequences of actions and outcomes to ensure continuity and coherence in decision-making; ii) Long-Term Summary Memory: A condensed knowledge base that captures and abstracts essential patterns, lessons, and strategic insights accumulated throughout the mission. By integrating these components, the system can perform context-aware prompt retrieval, leveraging both immediate context and historical understanding. This approach supports sustained strategic awareness and reasoning stability over extended deployments, improving reliability and adaptability in complex public safety operations. Overall, while DRL algorithms are state of the art for optimizing UASNETs, they reach their limits in emergency scenarios due to their slow training, complex application and gaps between simulation and reality, making them unsuitable for time-critical situations. DMs improve DRL by generating synthetic data and improving the adaptability of policies, but are still computationally intensive. In contrast, LLM optimizers that use ICL offer a promising alternative. They enable fast, training-free adaptation through natural language prompts, iterative refinement of the solution, and optimization that can be interpreted by humans - important advantages for emergency response. However, LLM-based approaches require further validation of their convergence properties and real-world reliability, suggesting a potential for hybrid systems that combine the robustness of DRL with the flexibility of LLM to meet urgent public safety needs. Table II Compare Optimization Approaches for Public Safety UAV and provide details on training time and inference latency.
@@ VII Most recent trends
This section highlights the future direction of LLM-assisted ICL and the ways to improve the performance and deployment.
@@ VII-1 Enriching Examples using Diffusion Models
Focus on enhancing the generalization capabilities of public safety UAV by addressing the scarcity of high-quality training data. Develop methods to generate realistic synthetic examples using DMs to augment limited emergency datasets. Further research should explore optimizing DM-based data generation to ensure the quality of synthetic examples for robust generalization.
@@ VII-3 Enhanced Robustness Against Adversarial Attacks
Investigate and develop advanced defense mechanisms to mitigate vulnerabilities such as jailbreaking attacks, which manipulate prompts to bypass built-in safety mechanisms and cause LLMs to generate malicious content. Address both white-box and black-box categories of adversarial attacks to ensure the reliability of LLM-assisted ICL in mission-critical scenarios.
@@ VII-4 Extension to Public Safety UAV Swarms
Extend the application of LLM-assisted ICL to public safety UAV swarms, focusing on facilitating coordination among multiple UAVs. Develop collaborative reasoning strategies that allow UAVs to share contextual information and jointly adapt to dynamic environments. Develop methods that enable cooperative task execution in the swarm, leveraging the contextual understanding and adaptive reasoning capabilities of LLMs to improve overall swarm performance and mission success. Such an extension can support distributed decision-making, enhance resilience to single-node failure, and enable collective intelligence across heterogeneous UAV platforms.
@@ VII-5 Real-World Validation and Transferability
Conduct extensive field trials to evaluate the performance of LLM-assisted ICL in real-world public safety applications. These trials should assess latency, energy efficiency, communication reliability, and overall mission success under practical operating constraints. Explore the transferability of LLM-assisted ICL to other UAV applications, such as agriculture and infrastructure inspection, by repurposing task descriptions and prompts. Our future work will also focus on quantifying how well the proposed framework generalizes across domains and identifying the degree of prompt adaptation required for new operational contexts.